OpenStudy (anonymous):

write the slope intercept equation for the line with the given slope that contains the given point. m=8/3.,(3,-4)

OpenStudy (anonymous):

put into the equation y = mx + b you have the x and y values given to you in the co-ordinates and b is the y-intercept that you're looking for. (-4) = (8/3)(3) + b (-4) = 8 + b (-4) - 8 = b b = -12 equation of the line: y = (8/3)x - 12

OpenStudy (anonymous):

thx =)